In the ABO blood groups, there are A, B, O, and AB blood types. They are due to three alleles, IA, IB, and IO, with IAIA and IAIO having blood group A; with IBIB and IBIO having blood group B; with IAIB having blood group AB; and with IOIO having blood group O. Among many aboriginal American Indian tribes, the allele frequency of IB in the ABO blood groups is extremely low. In one sample of 600 Papago Indians from Arizona, there were 37 A and 563 O blood types (no B and no AB). What are the best estimates of the allele frequencies of IA, IB and IO in this population, and what are the expected genotype frequencies assuming HWE?
Answer #1

In Hardy Weinberg equilibrium p²+q²+2pq = 1

Here O blood group is present in 567 individuals out of 600 and we know that O blood is due to homozygous condion that is IOIO.

So, frequency of IO, that is, p in population will be

p²= 567/600= 0.945

p= 0.972

We also know that p+q=1

q= 1- 0.972= 0.028

Therefore frequency of IA in population will be 0.028.

Now the genotypic frequency

IAIA homozygous will be (0.028)²= 0.00784

IOIO will be 0.945

IAIO will be 2pq= 0.0544

And there is no presence of IB allele so its frequency will be zero.